A total of 320 tickets were sold for the school play. They were either adult tickets or student tickets. The number of student tickets sold was three times the number of adult tickets sold. How many adult tickets were sold
Answer by ikleyn(52793)

.

You write an equation as you read the problem


    a + 3a = 320    ("a" is the number of adult tickets)


Then you simplify it and find the answer


    4a     = 320

     a     = 320/4 = 80.


ANSWER.  80 adult tickets.

Solved.